OpenStudy (donnie1999):

HELP ME!!! What is the equation of the directrix of a parabola defined by the equation y = x^2 − 2x − 3?

OpenStudy (donnie1999):

@aloud

OpenStudy (anonymous):

plug it in the desmoes

OpenStudy (donnie1999):

IDK wat a directrix is

OpenStudy (anonymous):

If the parabola is rotated so that its vertex is (h,k) and its axis of symmetry is parallel to the x-axis, it has an equation of (y - k)^2 = 4p (x - h), where the focus is (h + p, k) and the directrix is x = h - p.

OpenStudy (donnie1999):

So would it be y=4

OpenStudy (donnie1999):

@aloud would the answer be y=4

OpenStudy (anonymous):

y=-4 x=1

OpenStudy (donnie1999):

so that a yes lol

OpenStudy (anonymous):

yes